OpenStudy (anonymous):

how to multiply 5/6 times 24

OpenStudy (anonymous):

think of 24 as a fraction \[\frac{ 5 }{ 6 }*\frac{ 24 }{ 1 }\]

OpenStudy (anonymous):

When fractions are multiplied together the numerators are multiplied into a new numerator and the denominators are multiplied into a new denominator

OpenStudy (anonymous):

\[\frac{ a }{ b }*\frac{ c }{ d }=\frac{ a*c }{ b*d }\]

OpenStudy (anonymous):

thanks. (:

OpenStudy (anonymous):

yw